# Budgeted Spatial Data Acquisition: When Coverage and Connectivity Matter

### v2

@@ Abstract
Data is undoubtedly becoming a commodity like oil, land, and labor in the 21st century. Although there have been many successful marketplaces for data trading, the existing data marketplaces lack consideration of the case where buyers want to acquire a collection of datasets (instead of one), and the overall spatial coverage and connectivity matter. In this paper, we take the first attempt to formulate this problem as Budgeted Maximum Coverage with Connectivity Constraint (BMCC), which aims to acquire a dataset collection with the maximum spatial coverage under a limited budget while maintaining spatial connectivity. To solve the problem, we propose two approximate algorithms with detailed theoretical guarantees and time complexity analysis, followed by two acceleration strategies to further improve the efficiency of the algorithm. Experiments are conducted on five real-world spatial dataset collections to verify the efficiency and effectiveness of our algorithms.
@@ 1 Introduction
Data is often referred to as the new oil of the digital economy, representing a highly valuable and untapped asset [1, 2]. To fully realize the potential of data, various data marketplace platforms [3, 4, 5, 6, 7, 8] have emerged and offer a diverse range of data types for trade. For instance, Xignite [4] sells financial data, Gnip [5] sells data from social media, and Datarade [3] sells multiple categories of data, including commercial, academic, and geospatial data, etc. Among them, geospatial data has become increasingly important. According to a research report by Verified Market [9], the global geospatial analytics market was valued at \$66 billion in 2021 and is expected to reach \$209 billion by the end of 2030.
@@
Driven by the enormous economic profit, the geospatial data marketplaces, where buyers acquire or purchase data provided by sellers, are experiencing rapid growth. For example, the EU-funded OpertusMundi project [10] develops the Topio [11] to sell geospatial datasets of Europe, the UP42 [12] sells a wide range of earth observation dataset, and AggData [6] sells various location datasets such as restaurants and stations.
@@
The Problem We Study - Collection Acquisition. In this paper, we take the first attempt to formulate the data collection acquisition problem as the Budgeted Maximum Coverage with Connectivity Constraint (BMCC), which aims to find a set of datasets with the maximum spatial coverage under a limited budget while maintaining spatial connectivity. We solve this problem by casting it into a graph problem. Specifically, we construct a graph to model dataset relationship, where each node represents a dataset, and each edge represents that two datasets are connected. The objective of BMCC is to find a set of nodes (the total price within the budget) that induces a connected subgraph to maximize spatial coverage.
We prove BMCC to be NP-hard and propose two approximate algorithms with theoretical guarantees, namely Dual-Search Algorithm (DSA) and Dual-Path Search Algorithm (DPSA), followed by two acceleration strategies. The basic idea of DSA is to iteratively pick the dataset, which brings the maximum marginal gain w.r.t. the spatial coverage while maintaining the spatial connectivity. However, the theoretical analysis shows that the approximation ratio of DSA gradually decreases as the budget increases. To address this, we propose DPSA, which iteratively selects paths (i.e., a sequence of nodes connected by edges) with the common node and the maximum marginal gain from the dataset graph. The theoretical analysis shows its better approximation in scenarios involving a larger budget. Our contributions are summarized below:
@@
  • To further improve the approximation ratio with large input budgets, we propose the DPSA with detailed theoretical guarantees and time complexity analysis. Furthermore, we propose two acceleration strategies to enhance the efficiency of the DPSA significantly.
@@ 2 Related Work
The literature on data marketplace can be broadly divided into three classes: (i) data marketplace model [27, 20, 2, 28, 19], (ii) data pricing [29, 30, 31, 32, 33, 34], and (iii) data acquisition [23, 21, 35, 36, 37]. We will first provide an overview of each of these classes. In the subsequent sections, we will introduce several combinatorial optimization problems closely related to BMCC, such as the influence maximization (IM) problem [38, 39, 40, 41, 42, 43, 44, 45], dominating set (DS) problem [46, 47, 48, 49, 50, 51, 52], and maximum coverage problem (MCP) [18, 53, 54, 55, 56, 57, 58, 59].
@@
Data Marketplace Model.  A data marketplace model [27, 20, 2, 28, 60] consists of buyers who want to buy datasets, sellers who want to share datasets in exchange for a reward, and an arbiter who satisfies the buyers’ requests with the seller’s supplied datasets. For instance, Agarwal et al. [61] propose a data model that solves the allocation and payment problem in a static scenario with one buyer and multiple sellers. In [62], the authors propose an end-to-end market design that considers buyers and sellers arriving in a streaming fashion. Additionally, Fernandez et al. [2] present their vision for the design of platforms and design the market rules to govern the interactions between sellers, buyers, and an arbiter.
@@
Data Pricing.  Data pricing [33, 32, 29, 63, 30, 31, 64, 65, 66, 67] has been widely studied by the communities of computer science and economics (see [68, 69] for a survey). For instance, for query-based pricing, Koutris et al. [30] propose the first pricing system named QueryMarket, which supports the efficient computation of prices for multiple queries using integer linear program solvers. Deep et al. [70] propose history-aware pricing that supports buyer charges according to current and past queries, which can support pricing in real time. For machine learning-based (ML) pricing, Chen et al. [33] first formally describe the desired properties of the model-based pricing framework and provide algorithmic solutions. In addition, there are many other types of pricing mechanisms, such as feature-based pricing [71, 72], usage-based pricing [73], etc. Note that our study is orthogonal to the data pricing studies. Our algorithm proposed in this paper can work with any pricing mechanism. For simplicity, we just use the usage-based pricing strategy [73] for illustration.
@@
Data Acquisition.  Data acquisition is another key factor in helping data buyers obtain high-quality data from the marketplace. For example, Li et al. [21] propose a middleware called DANCE, which aims to acquire a target table from the data marketplace with a high correlation between its attributes and query table. Chen et al. [37] consider the problem of obtaining data from multiple data providers in order to improve the accuracy of linear statistical estimators. Li et al. [35] investigate the problem of data acquisition by a specific query predicate (e.g., ten images with label = ‘dog’) and budget for improving the performance of ML models. Similarly, Asudeh et al. [23] develop a query-answering framework to acquire an integrating dataset that meets user-provided query predicate and distribution requirements (e.g., 100 cancer datasets with at least 20% black patients). Despite notable progress being made in this direction, existing studies primarily concentrate on tabular data, relying on tabular attributes or metadata information to acquire datasets that fulfill query requests. However, these approaches are not directly applicable to the acquisition of spatial datasets that aim to maximize coverage while ensuring connectivity.
@@
Combinatorial Optimization Problem.  A combinatorial optimization problem involves identifying the optimal combination of variables that maximizes a given objective value from a set of options, while adhering to specific constraints. For example, the budgeted influence maximization (BIM) problem [40, 74, 75, 76] aims to select a limited number of seed nodes such that the total cost of nodes is less than the predefined budget, and the influence spread is maximized. The budgeted connected dominating set (BCDS) problem [49, 48, 50, 51, 46] is to find $k$ nodes ($k$ is the budget) that induce a connected subgraph, and the number of nodes dominated by the $k$ nodes is maximized. The goal of the maximum coverage problem (MCP) is to select a limited number of sets, such that the number of the covered elements in these chosen sets is maximal. They focus on selecting the combination of sets in order to maximize the objective score, while the definitions of objective scores depend on specific problem contexts. For instance, the objective scores in the BIM, BCDS, and MCP correspond to the maximization of influencers, the number of dominated nodes, and the number of the set union, respectively. Although these variants appear to have similar constraints, such as budget, connectivity, etc., they face different challenges due to different objective scores, which is also the reason why we need to design specialized algorithms for our proposed BMCC problem.
@@
Difference between BMCC, BIM, and BCDS.  For the BIM, the challenge is how to measure the influence spread of a node since the calculation of influence spread from the seed set was proven to be an NP-complete problem [42]. For the BCDS, the challenge is how to maintain a connected dominating subset such that the number of nodes dominated by the subset is maximized [46]. However, our problem has a unique challenge: maximizing the union of elements covered by the selected subset under the connectivity and budget constraints. Therefore, the existing approaches to the BIM and BCDS problems cannot address our problem.
@@
Difference between BMCC and MCP.  Our problem can be viewed as a variant of the MCP. Hochbaum and Pathria [18] showed that the MCP is NP-hard and devised a simple greedy algorithm with an approximation factor of 1-$1/e$. Moreover, there are many variants that have been extensively studied, such as the budgeted MCP (BMCP) [77], the weighted MCP (WMCP) [78], the connected MCP (CMCP) [79], and the generalized MCP (GMCP) [80]. These variants are optimization problems that share the common objective of selecting sets to maximize the total weight of elements in the union set (the weight of each element in the MCP is 1). However, they differ in the constraints imposed on the solution. With different constraints, existing studies on these problems face different challenges, and effective solutions require problem-specific designs. It is unrealistic to have a general algorithm that fits all MCP variants above.
For the MCP and its variants, the extended algorithms based on the greedy strategy have been shown to have the best-possible approximation in polynomial time [81]. For example, for the BMCP, the greedy algorithm gives a 1/2(1-1/e)-approximation for any budget and a (1-1/e)-approximation when the budget can buy at least three sets [77]. For the CMCP, the greedy algorithm gives a 1/($c\cdot r_{OPT}$)-approximation (c = $\frac{2e-1}{e-1}\approx$2.58 and $r_{OPT}$ is the radius of optimal solution) [79]. Thus, we first try an extension of the greedy algorithm to solve our problem and perform a theoretical approximation analysis in Section 5.1. However, the theoretical analysis shows that the approximation ratio of DSA gradually decreases as the budget increases. To address this, we propose an improved algorithm in Section 5.2 to produce better approximation ratios when large budgets are involved.
@@ 3 Problem Formulation
In this section, we introduce the data modeling and formalize the problem. Frequently used notations are shown in Table 1.
@@ 3.1 Data Modeling
To quantify the coverage of a spatial dataset, we partition the space into a grid of uniform cells [82, 83], since the individual points have no length or width. This approach helps in measuring the dataset’s coverage more effectively. The definitions are as follows.
Definition 3.
[…]
@@
Fig. 2(a) first shows an example of a raw spatial dataset $D$ in space. Each point of $D$ can be mapped to the cell of the grid, as shown in Fig. 2(b). Each cell can be uniquely represented by an integer, e.g., for the bottom left cell, its ID is 00, which is transformed from its coordinates (0, 0). Thus, the spatial dataset $D$ can be represented as a finite set $S_{D}$ consisting of a sequence of cell IDs, as shown in Fig. 2(c). The spatial coverage of $S_{D}$ is the number of IDs in the set, which is 5.
@@
$\mathcal{D}$ the data marketplace
@@
$dist(S_{D^{\prime}},S_{D})$ the distance between $S_{D^{\prime}}$ and $S_{D}$
@@
(Cell-based Dataset Distance) Given two cell-based datasets $S_{D}$ and $S_{D^{\prime}}$, their distance $dist(S_{D},S_{D^{\prime}})$ is the minimum distance between their cells and is defined as:
$\small dist(S_{D},S_{D^{\prime}})=\min_{c\in S_{D},c^{\prime}\in S_{D^{\prime}}}\>||c,c^{\prime}||_{2}$ (1)
@@
In addition, to evaluate whether a set of spatial datasets satisfies the spatial connectivity, it is essential to establish criteria defining the relationships between pairs of datasets. Thus, we introduce the concepts of direct and indirect connected relations between two spatial datasets.
@@
(Directly Connected Relation) Given a distance threshold $\delta$, we say two cell-based datasets $S_{D}$ and $S_{D^{\prime}}$ are directly connected if $dist(S_{D},S_{D^{\prime}})\leq\bm{\delta}$.
Definition 7.
(Indirectly Connected Relation) Two cell-based datasets $S_{D}$ and $S_{D^{\prime}}$ are indirectly connected if there exists one or multiple ordered datasets $\{S_{D_{1}},\dots,S_{D_{i}}\}(i\geq 1)$ such that each pair of adjacent datasets (e.g., ($S_{D},S_{D_{1}}$), …, ($S_{D_{i}},S_{D^{\prime}}$)) are directly connected relation.
Definition 8.
(Spatial Connectivity) Given a collection of cell-based datasets $\mathcal{S}_{\mathcal{D}}=\{S_{D_{1}},\ldots,S_{D_{|\mathcal{D}|}}\}$ constructed from $\mathcal{D}$ and a grid $\mathcal{C}$, the collection $\mathcal{S}_{\mathcal{D}}$ satisfies the spatial connectivity if and only if any pair of datasets $S_{D_{i}}$ and $S_{D_{j}}$ in $\mathcal{S}_{\mathcal{D}}(i\neq j)$ are directly or indirectly connected relation.
@@
((Connected) Spatial Dataset Graph) Given a set of cell-based datasets $\mathcal{S}_{\mathcal{D}}=\{S_{D_{1}},\ldots,S_{D_{|\mathcal{D}|}}\}$ constructed from $\mathcal{D}$ and a grid $\mathcal{C}$, a pricing function $p$, and a distance threshold $\delta$, we can construct an undirected graph $\mathcal{G}_{\mathcal{S}_{\mathcal{D}},\delta}=(\mathcal{V},\mathcal{E})$, where each $v_{i}\in\mathcal{V}$ denotes a cell-based dataset $S_{D_{i}}\in\mathcal{S}_{\mathcal{D}}$ with a fixed price $p(S_{D_{i}})$ and there is an edge between $v_{i}$ and $v_{j}\in\mathcal{V}$ $(i\neq j)$, if and only if $S_{D_{i}}$ and $S_{D_{j}}$ are directly connected relation. The graph $\mathcal{G}_{\mathcal{S}_{\mathcal{D}},\delta}$ is a connected spatial dataset graph if the collection $\mathcal{S}_{\mathcal{D}}$ satisfies the spatial connectivity.

### v3

@@ Abstract
Data is undoubtedly becoming a commodity like oil, land, and labour in the 21st century. Although there have been many successful marketplaces for data trading, the existing data marketplaces lack consideration of the case where buyers want to acquire a collection of datasets (instead of one), and the overall spatial coverage and connectivity matter. In this paper, we make the first attempt to formulate this problem as Budgeted Maximum Coverage with Connectivity Constraint (BMCC), which aims to acquire a dataset collection with the maximum spatial coverage under a limited budget while maintaining spatial connectivity. To solve the problem, we propose two approximate algorithms with detailed theoretical guarantees and time complexity analysis, followed by two acceleration strategies to further improve the efficiency of the algorithm. Experiments are conducted on five real-world spatial dataset collections to verify the efficiency and effectiveness of our algorithms.
@@ 1 Introduction
Data is often referred to as the new oil of the digital economy, representing a highly valuable and untapped asset [1, 2]. To fully realize the potential of data, various data marketplace platforms have emerged and offer a diverse range of data types for trade. For instance, Xignite sells financial data, Gnip sells data from social media, and Datarade sells multiple categories of data, including commercial, academic, and spatial data, etc. Among them, spatial data has become increasingly important. According to a research report by Verified Market, the global spatial analytics market was valued at \$66 billion in 2021 and is expected to reach \$209 billion by the end of 2030.
@@
Driven by the enormous economic profit, the geospatial data marketplaces, where buyers acquire or purchase data provided by sellers, are experiencing rapid growth. For example, the EU-funded OpertusMundi project develops the Topio [3] to sell geospatial datasets of Europe, the UP42 sells a wide range of earth observation dataset, and AggData sells various location datasets such as restaurants and stations.
@@
The Problem We Study - Collection Acquisition. In this paper, we take the first attempt to formulate the data collection acquisition problem as the Budgeted Maximum Coverage with Connectivity Constraint (BMCC), which aims to find a set of datasets with the maximum spatial coverage under a limited budget while maintaining spatial connectivity. To solve the maximization problem under multiple constraints at the same time, we first transform it into a graph theory problem. Specifically, we construct a graph to model dataset relationships, where each node represents a dataset, and each edge represents that two datasets are connected. The objective of BMCC is to find a set of nodes (the total price within the budget) that induces a connected subgraph to maximize spatial coverage.
The difficulty of the BMCC problem arises from its NP-hard nature, which refers to a problem for which no known polynomial-time algorithm can find a solution. We formally prove that BMCC is NP-hard by using a polynomial-time reduction from the classic maximum coverage problem (MCP) [9] in Section 4. Inspired by the greedy algorithm proposed by Fiege [10], which is the best possible polynomial time approximation algorithm for the maximum coverage problem and its variants. Thus, we propose two heuristic algorithms based on greedy strategy to solve the BMCC problem. The intuition behind these algorithms is to iteratively select the best available candidate dataset that maximizes the coverage, while considering the constraints imposed by the problem.
Specifically, we propose two approximate algorithms with theoretical guarantees, namely Dual-Search Algorithm (DSA) and Dual-Path Search Algorithm (DPSA), followed by two acceleration strategies. The basic idea of DSA is to iteratively pick the dataset, which brings the maximum marginal gain w.r.t. the spatial coverage while maintaining the spatial connectivity. However, the theoretical analysis shows that the approximation ratio of DSA gradually decreases as the budget increases. To address this, we propose DPSA, which iteratively selects paths (i.e., a sequence of nodes connected by edges) with the common node and the maximum marginal gain from the dataset graph. The theoretical analysis shows its better approximation in scenarios involving a larger budget. Our contributions are summarized below:
@@
  • To further improve the approximation ratio with large input budgets, we propose DPSA with detailed theoretical guarantees and time complexity analysis. Furthermore, we propose two acceleration strategies to enhance the efficiency of DPSA significantly.
@@ 2.1 Data Marketplace
The literature on data marketplace can be broadly divided into three classes: (i) data marketplace model [19, 20, 2], (ii) data pricing [21, 22, 23, 24, 25, 26], and (iii) data acquisition [15, 13, 27]. We will provide an overview of each of these classes.
@@
Data Marketplace Model.  A data marketplace model [28, 12, 2, 29, 30, 19, 20] consists of buyers who want to buy datasets, sellers who want to share datasets in exchange for a reward, and an arbiter who satisfies the buyers’ requests with the seller’s supplied datasets. For instance, Agarwal et al. [19] propose a data model that solves the allocation and payment problem in a static scenario with one buyer and multiple sellers. In [20], the authors propose an end-to-end market design that considers buyers and sellers arriving in a streaming fashion. Additionally, Fernandez et al. [2] present their vision for the design of platforms and design the market rules to govern the interactions between sellers, buyers, and an arbiter.
@@
Data Pricing.  Data pricing [25, 24, 21, 22, 23, 31, 32] has been widely studied by the communities of computer science and economics (see [33, 34] for a survey). For instance, for query-based pricing [35, 22], Koutris et al. [22] propose the first pricing system named QueryMarket, which supports the efficient computation of prices for multiple queries using integer linear program solvers. Deep et al. [36] propose history-aware pricing that supports buyer charges according to current and past queries, which can support pricing in real time. For machine learning-based (ML) pricing [37], Chen et al. [25] first formally describe the desired properties of the model-based pricing framework and provide algorithmic solutions. In addition, there are many other types of pricing mechanisms, such as feature-based pricing [38, 39], usage-based pricing [40], etc. Note that our study is orthogonal to the data pricing studies. Our algorithm proposed in this paper can work with any pricing mechanism. For simplicity, we just use the usage-based pricing strategy [40] for illustration.
@@
Data Acquisition.  Data acquisition is another key factor in helping data buyers obtain high-quality data from the marketplace. For example, Li et al. [13] propose a middleware called DANCE, which aims to acquire a target table from the data marketplace with a high correlation between its attributes and query table. Chen et al. [41] consider the problem of obtaining data from multiple data providers in order to improve the accuracy of linear statistical estimators. Li et al. [27] investigate the problem of data acquisition by a specific query predicate (e.g., ten images with label = ‘dog’) and budget for improving the performance of ML models. Similarly, Asudeh et al. [15] develop a query-answering framework to acquire an integrating dataset that meets user-provided query predicate and distribution requirements (e.g., 100 cancer datasets with at least 20% black patients).
Despite notable progress being made in this direction, existing studies primarily concentrate on tabular data, which typically refers to datasets organized in tables with rows and columns. In contrast to tabular data, spatial data has spatial attributes associated with each data point (e.g., latitude and longitude). Although spatial data can be converted into tabular form to some extent, this representation may not adequately reflect its spatial characteristics and complexity. Thus, these approaches for acquiring tabular datasets are not directly applicable to the acquisition of spatial datasets. This is because tabular data acquisition typically relies on attributes or metadata to query for relevant datasets (e.g., filtering rows based on column values), while spatial data acquisition involves spatial relationships and geometry, such as proximity or spatial overlap. The complexity of spatial relationships, combined with the need to consider coverage and connectivity, requires more specialized methods that consider not only attribute similarity but also spatial features and their relative positions in space.
2.2 Combinatorial Optimization Problem
A combinatorial optimization problem involves identifying the optimal combination of variables that maximizes a given objective value from a set of options, while adhering to specific constraints. In the subsequent sections, we will discuss difference between several classic combinatorial optimization problems that have similar constraints and objectives with BMCC, such as the budgeted influence maximization (BIM) problem [42, 43, 44, 45, 46, 47, 48, 49], budgeted connected dominating set (BCDS) problem [50, 51, 52, 53, 54, 55, 56], and maximum coverage problem (MCP) [9, 57, 58, 59, 60, 61, 62], and illustrate the technical innovations of our proposed method.
@@ 2.2 Combinatorial Optimization Problem
Difference between BMCC, BIM, and BCDS.  The budgeted influence maximization (BIM) problem [44, 63, 64, 65] aims to select a limited number of seed nodes such that the total cost of nodes is less than the predefined budget, and the influence spread is maximized. While the budgeted connected dominating set (BCDS) problem [53, 52, 54, 55, 50] is to find $k$ nodes ($k$ is the budget) that induce a connected subgraph, and the number of nodes dominated by the $k$ nodes is maximized. They focus on selecting the combination of sets in order to maximize the objective score, while the definitions of objective scores depend on specific problem contexts. For instance, the objective scores in BIM, BCDS, and BMCC correspond to the maximization of influencers, the number of dominated nodes, and the spatial coverage, respectively.
@@
For the BIM, the challenge is how to measure the influence spread of a node since the calculation of influence spread from the seed set was proven to be an NP-complete problem [46]. For the BCDS, the challenge is how to maintain a connected dominating subset such that the number of nodes dominated by the subset is maximized [50]. Although these optimization problems appear to have similar constraints, such as budget, connectivity, etc., our problem has a unique challenge: maximizing the union of elements covered by the selected subset under the connectivity and budget constraints. Therefore, the existing approaches to the BIM and BCDS problems cannot address our problem.
@@
Difference between BMCC and MCP.  The goal of MCP is to select a limited number of sets, such that the number of the covered elements in these chosen sets is maximal. It is noted that our problem can be viewed as a variant of MCP. Hochbaum and Pathria [9] showed that MCP is NP-hard and devised a greedy algorithm with an approximation factor of 1-$1/e$. Moreover, there are many variants that have been extensively studied, such as budgeted MCP (BMCP) [66], weighted MCP (WMCP) [67], connected MCP (CMCP) [68], and generalized MCP (GMCP) [69]. These variants share a common goal, which is to select sets that maximize the total weight of the elements in the union set. However, they differ in the constraints imposed on the solution.
For these combinatorial optimization problems like the MCP and its variants, heuristic algorithms can be employed to find solutions, such as greedy algorithm [59], ant colony optimization [70], et al. However, algorithms based on greedy strategies have been shown to have the best possible approximation in polynomial time [10]. A straightforward adaptation based on the greedy strategy is to first find the dataset with maximum coverage and iteratively choose datasets with the maximum coverage increment that satisfies the spatial connectivity and budget constraints. Since this straightforward approach requires evaluating the spatial connectivity of the resulting dataset at each iteration, this approach can be expensive when the number of datasets is large and lacks any formal theoretical guarantees regarding the approximation ratio.
Novelty of Our Methods. The novelty of our methods lies in the proposed greedy and acceleration strategies with theoretical approximation guarantees. Specifically, we first design a DSA algorithm that incorporates an index-based acceleration strategy to handle the evaluation of spatial connectivity more efficiently, thus reducing computational overhead. However, the theoretical analysis shows that the approximation ratio of DSA gradually decreases as the budget increases. To overcome the limitations of DSA, the DPSA algorithm is introduced based on the designed spatial dataset graph. It iteratively selects paths with common nodes and the maximum marginal gain from the dataset graph. The DPSA algorithm demonstrates better approximation performance, especially for larger budgets. In addition to the algorithms, we also propose two acceleration strategies in Section 5.2.2 to enhance the efficiency of these algorithms.
@@ 3 Problem Formulation
In this section, we introduce the data modeling and formalize the problem. A summary of frequently used notations is provided in Table 1.
@@ 3.1 Data Modeling
To quantify the coverage of a spatial dataset, we partition the space into a grid of uniform cells [71, 72], since the individual points have no length or width. This approach provides a more effective means of measuring dataset coverage. Specifically, given an integer $\theta$, we partition the entire space containing all spatial datasets into a grid $\mathcal{C}_{\theta}$ of $2^{\theta}\times 2^{\theta}$ uniform cells, where $\theta$ is called the resolution.
Each cell, denoted as $c$, denotes a unit space in the grid $\mathcal{C}_{\theta}$. Its coordinates $(x,y)$ can be converted into a non-negative integer to uniquely indicate $c$ using the z-order curve method [73, 74], denoted as $z(x,y)=c$. This transformation yields consecutive IDs in the range $[0,2^{\theta}\times 2^{\theta}-1]$. In this way, a spatial dataset $D$ can be represented as a sorted integer set consisting of a sequence of cell IDs, referred to as a cell-based dataset. The spatial coverage of a dataset is the number of cell IDs in its corresponding cell-based dataset. The formal definition of cell-based dataset and spatial coverage is as follows.
@@
Fig. 2(a) first shows an example of a raw spatial dataset $D$ in space. Each point of $D$ can be mapped to the cell of the grid, as shown in Fig. 2(b). Each cell can be uniquely represented by an integer, e.g., for the bottom left cell, its ID is 00, which is transformed from its coordinates (0, 0). Thus, the spatial dataset $D$ can be represented as a finite set $S_{D}$ consisting of a sequence of cell IDs, i.e., $S_{D}=\{3,6,9,11,12\}$, as shown in Fig. 2(c). The spatial coverage of $S_{D}$ is the number of IDs in the set, which is 5.
@@
$D$ a spatial dataset
$\mathcal{D}$ a spatial data marketplace
@@
$dist(S_{D_{i}},S_{D_{j}})$ the distance between $S_{D_{i}}$ and $S_{D_{j}}$
@@
(Cell-based Dataset Distance) Given two cell-based datasets $S_{D_{i}}$ and $S_{D_{j}}$, their distance $dist(S_{D_{i}},S_{D_{j}})$ is the minimum distance between their cells and is defined as:
$\small dist(S_{D_{i}},S_{D_{j}})=\min_{c\in S_{D_{i}},c^{\prime}\in S_{D_{j}}}\>||c,c^{\prime}||_{2}$ (1)
@@
In addition, to evaluate whether a set of spatial datasets satisfies the spatial connectivity, it is essential to establish criteria defining the relationships between pairs of datasets. Thus, we introduce the concepts for defining relationships between datasets. Thus, we introduce the concept of spatial connectivity, which encompasses both direct and indirect connections.
@@
(Spatial Connectivity) Given a collection of cell-based datasets $\mathcal{S}_{\mathcal{D}}=\{S_{D_{1}},\ldots,S_{D_{|\mathcal{D}|}}\}$ constructed from $\mathcal{D}$ and a grid $\mathcal{C}_{\theta}$, the collection $\mathcal{S}_{\mathcal{D}}$ satisfies the spatial connectivity if, for a given distance threshold $\delta$, any pair of datasets $S_{D_{i}}\in\mathcal{S}_{\mathcal{D}}$ and $S_{D_{j}}\in\mathcal{S}_{\mathcal{D}}$ $(i\neq j)$ are either directly connected (i.e., $dist(S_{D_{i}},S_{D_{j}})\leq\bm{\delta}$) or indirectly connected through a sequence of ordered datasets $\{S_{D_{m}},\dots,S_{D_{n}}\}$, where each pair of adjacent datasets (i.e., ($S_{D_{i}},S_{D_{m}}$), …, ($S_{D_{n}},S_{D_{j}}$)) directly connected.
@@
((Connected) Spatial Dataset Graph) Given a set of cell-based datasets $\mathcal{S}_{\mathcal{D}}=\{S_{D_{1}},\ldots,S_{D_{|\mathcal{D}|}}\}$ constructed from $\mathcal{D}$ and a grid $\mathcal{C}_{\theta}$, a pricing function $p$, and a distance threshold $\delta$, we can construct an undirected graph $\mathcal{G}_{\mathcal{S}_{\mathcal{D}},\delta}=(\mathcal{V},\mathcal{E})$, where each $v_{i}\in\mathcal{V}$ denotes a cell-based dataset $S_{D_{i}}\in\mathcal{S}_{\mathcal{D}}$ with a fixed price $p(S_{D_{i}})$ and there is an edge between $v_{i}$ and $v_{j}\in\mathcal{V}$ $(i\neq j)$, if and only if $S_{D_{i}}$ and $S_{D_{j}}$ are directly connected relation. The graph $\mathcal{G}_{\mathcal{S}_{\mathcal{D}},\delta}$ is a connected spatial dataset graph if the collection $\mathcal{S}_{\mathcal{D}}$ satisfies the spatial connectivity.